OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)]

OpenStudy (anonymous):

g(-2)=6(-2)-1 = -12-1=-13 f(g(-2))= f(-13)= -2(-13) -1 = 26 -1=25

OpenStudy (anonymous):

so the answer is -13

OpenStudy (anonymous):

The answer is 25